OpenStudy (anonymous):

Given the trinomial, what is the value of the coefficient 'B' in the factored form? 2x2 + 14xy - 36y2 = 2(x + By)(x - 2y) B= ____________

OpenStudy (tkhunny):

Have you considered multiplying the binomials and seeing what you get?

OpenStudy (anonymous):

I don't know how and I'm really in a hurry to get this assignment done. I have to be 80% complete with the course by Monday and Im only at 42%right now

OpenStudy (anonymous):

Can you explain how?

OpenStudy (tkhunny):

I don't believe you cannot multiply two binomials. Give it a go. (x + By)(x - 2y) = ??

OpenStudy (anonymous):

x^2+Bxy-2By^2. Is that the correct answer?

OpenStudy (tkhunny):

You missed "-2xy"

OpenStudy (anonymous):

So x^2+Bxy-2xy-2By^2. Is that better?

OpenStudy (tkhunny):

Okay, now somehow, 2(x2 + 7xy - 18y2) = 2(x^2+Bxy-2xy-2By^2) What value for B does that?

OpenStudy (anonymous):

@LearningIsAwesome

OpenStudy (anonymous):

@Sheraz12345

OpenStudy (tkhunny):

You didn't answ2er my question. Look at the y^2 term. What value for B makes those the same?

OpenStudy (tkhunny):

Considering ONLY the y^2 terms: - 18y^2 = -2By^2 What value of B will do that?

OpenStudy (anonymous):

2

OpenStudy (tkhunny):

Let's see, \(-2(2) = -4 \ne -18\) You're just guessing. \(2B = 18\) - Solve for B.

OpenStudy (anonymous):

B=9?

OpenStudy (tkhunny):

Very good. Do that again with the xy terms.

OpenStudy (tkhunny):

Well? 7xy = Bxy-2xy What value of B will do that?